40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 18 02:42:28 2014, Jenne, Update, LSC, AO path partly engaged Zoom_TRXTRY4_EngagingAO_1095059162.png0020.pdf
    Reply  Thu Sep 18 17:44:55 2014, Jenne, Update, LSC, Old AO cable pulled 
       Reply  Tue Sep 30 17:26:18 2014, ericq, Update, LSC, New AO cable in place 
    Reply  Fri Sep 19 04:05:05 2014, ericq, Update, LSC, AO path partly engaged TFSR785_19-09-2014_020555.pdfTFSR785_19-09-2014_033920.pdf
       Reply  Fri Sep 19 13:12:07 2014, Jenne, Update, LSC, AO path glitches 
Message ID: 10520     Entry time: Fri Sep 19 04:05:05 2014     In reply to: 10516     Reply to this: 10521
Author: ericq 
Type: Update 
Category: LSC 
Subject: AO path partly engaged 

More AO efforts. No huge news. 

Came at AO from each side. For each sign, I lost lock just a few dB from the AO portion of the loop crossing unity gain. Both attempts were about arm powers of 1, which should correspond to ~300pm CARM offset, which I have simulated the crossover as possible with my current loop models (including latest MC loop). The gain steps were usually 6dB in between measurements. 

Positive polarity on CM board screen:

I made it to +5 dB of the last plot here, but the 6th broke it open. Gains on CM In2, CM AO, and MC In2 were -6, -4, -2 on that last, lock breaking, step. 

TFSR785_19-09-2014_020555.pdf

Negative polarity on CM board screen:

Lost it just 2dB above the last trace. Gains were -6, +1, -2 (So, overall 5dB higher than the other polarization)

TFSR785_19-09-2014_033920.pdf

Many things happened in between these two lock stretches, but I'm not sure what may or may not have affected things. They include:

  • Jenne mentioned PRMI being fussy earlier in the evening. I adjusted REFL33 and POP22 angles during a PRMI lock, while CARM was held away with ALS. My simulations suggest that there are small changes to the 3F sensing when the arms are totally absent, but doing it at a finite CARM offset is closer to where we want it, it seems. 
  • I tried using REFL165Q for MICH, since my simulations suggest a better MICH/PRCL angle, which would stave off cross couplings. Lined up excitations, etc., but no luck. 
  • I measured the PRMI loops
    • found PRCL to have ~200Hz UGF, 8dB gain peaking. Maybe a little high, but didn't seem terrible. 
    • MICH had UGF of around 20Hz, with the FM gain at 0.8. By the shape of the phase bubble, the loop seems designed for higher bandwidth. I raised the gain to 2.5 for a 70ishHz UGF, and called in FMs 7 and 9 for additional triggered boosts. Things seemed to stay locked pretty well. 
  • Lower excitation amplitude the second time around, measuring the AO loop. Looking at the CM output spectra, you can see the excitation wailing away; I wanted to avoid it.

The location of the CARM resonance peak lines up with my simulation, which is good, but there appears to be less phase than expected... I tried making sure that we don't have any whitening uncompensated for, but it looked ok. All my AO path loop model contains is the CM board TF (measured and fitted), the IMC seen as an actuator(measured and fitted), and the REFLDC optical TF (simulated in MIST). Maybe the DC path of whatever diode this is coming from needs to be included...

Discontinuities / glitches could be seen in the CM board fast output when MC board gains were changed, which isn't so nice. Incidentally, I notice now that each lock loss corresponded to a step of AO gain on the CM board.

ELOG V3.1.3-